stranger2:

Geometry Which is the equation of a hyperbola with directrices at x = ±2 and foci at (4, 0) and (−4, 0)?

stranger2:

stranger2:

@dUdE

AZ:

I despise conic sections and I never enjoyed learning about it but I just brushed up on it since it seems like no one else wants to bite. I found this website to be particularly helpful in refreshing my knowledge and would advise you to take a look at it too. https://courses.lumenlearning.com/waymakercollegealgebra/chapter/equations-of-hyperbolas/ When the coordinates of the foci are \( (\pm c, ~0)\), the standard form of the equation of a hyperbola with center (0, 0) is \( \dfrac{{x}^{2}}{{a}^{2}}-\dfrac{{y}^{2}}{{b}^{2}}=1\) When the coordinates of the foci are \( (0,~ \pm c)\), the standard form of the equation of a hyperbola with center (0, 0) is \(\dfrac{{y}^{2}}{{a}^{2}}-\dfrac{{x}^{2}}{{b}^{2}}=1\) We're given the foci as (-4, 0) and (4, 0). Which standard form should we use? This should help you eliminate two answer choices almost immediately.

AZ:

Now, to finally narrow it down to the correct answer. The equation of directrices are \( x = \pm\dfrac{a^2}{c}\) We're given that the hyperbola has a directrices at \(x = \pm 2\) We know what c is because that's our foci. It's in the form \( ( \pm c, ~ 0)\) so our c is is 4. To find out what `a` and `b` is and get our equation, we need to first solve for `a` using \( x = \pm\dfrac{a^2}{c}\) since we know \(x = \pm 2\), we can set it up as \(\dfrac{a^2}{c} = 2\) plug in c = 4 and solve for a! To solve for b, you'll need to use the formula \( c^2 = a^2 + b^2\)

AZ:

Let me know what you get or if you get stuck anywhere along the way!

kevin565445432:

i need help

kevin5654444:

g-guys

stranger2:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Let me know what you get or if you get stuck anywhere along the way! \(\color{#0cbb34}{\text{End of Quote}}\) wowwww thankz ur such a genius

AZ:

It was my pleasure!